OpenStudy (idku):

let me try another Bernoulli DE (example 3)

OpenStudy (idku):

\(\large y'+e^{-x}y=y^7\)

OpenStudy (idku):

\(\large y'y^{-7}+e^{-x}y^{-6}=1\) \(\large\dfrac{-1}{6}v'+e^{-x}v=1\) \(\large v'-6e^{-x}v=-6\) \(\large v'e^{e^{6x}}-e^{e^{6x}}6e^{-x}v=-6e^{e^{6x}}\)

OpenStudy (idku):

\(\large ve^{e^{6x}}=\int -6e^{e^{6x}}dx\) well, I guess I am not yet able to integrate this, but \(\large y^{-6}=\dfrac{\int -6e^{e^{6x}}dx}{e^{e^{6x}}}\) \(\large y^{}=\dfrac{\left(e^{e^{6x}}\right)^6}{\left(\int -6e^{e^{6x}}dx\right)^6}\)

OpenStudy (idku):

Can someone give me a better example?

OpenStudy (idku):

\(\large y'-(1/x)y=y^4\)

OpenStudy (empty):

Two things! Slight mistake on exponents here: \(\large y^{}=\dfrac{\left(e^{e^{6x}}\right)^{1/6}}{\left(\int -6e^{e^{6x}}dx\right)^{1/6}}\) Also, if you really needed this integral, you could use the taylor series: \[\ e^{e^{6x}}=\sum_{n=0}^\infty \frac{e^{6nx}}{n!}\] Then integrate term by term, until you have a reasonable enough approximation for whatever application you want to use it for, cause like let's face it in the real world 10 digits of accuracy is not bad at all haha

OpenStudy (idku):

yes those are 1/6's

OpenStudy (idku):

I didn't know about that taylor series. Something to work on in my little math world.

OpenStudy (idku):

making a sandbox -:(

OpenStudy (idku):

I will do that taylor series at some other time, I want to digest the Bernoulli technique first.

OpenStudy (empty):

Oh you probably already know this power series! \[e^y=\sum_{n=0}^\infty \frac{y^n}{n!}\] \[y=e^{6x}\]

OpenStudy (idku):

oh just a sub, I am so ...don';t want to violate the site's policy.

OpenStudy (empty):

Haha hey at least it turned out being easier than you thought rather than it being harder than you thought! xD

OpenStudy (idku):

I should have figured that (after having taken calc 2)

OpenStudy (idku):

thanks.

OpenStudy (idku):

I will proceed to my next example I started to do then for now.....

OpenStudy (idku):

\(\large y'y^{-4}-(1/x)y^{-3}=1\) \(\large v'/(-3)-(1/x)v=1\) \(\large v'+(3/x)v=3\) \(\large v'e^{3\ln x}+e^{3\ln x}(3/x)v=e^{3\ln x}\) \(\large ve^{3\ln x}=e^{3\ln x}\) \(\large ve^{3\ln x}=\int x^3 dx \) \(\large vx^3=(1/4)x^4+C \) \(\large v=\dfrac{\frac{1}{4}x^4+C}{x^3}\) \(\large y^{-3}=\dfrac{\frac{1}{4}x^4+C}{x^3}\) \(\large y=\dfrac{(\frac{1}{4}x^4+C)^{3}}{x^9}\)

OpenStudy (idku):

I left out the integral in line 5

OpenStudy (idku):

but I fixed it in line 6

OpenStudy (empty):

Are you verifying your solutions by plugging them back into the original differential equation?